find the value(s) of x for which f(x)=g(x).
f(x)=square root (3x)+1
g(x)=x+1
: find the value(s) of x for which f(x)=g(x).
f(x)=square root (3x)+1
g(x)=x+1

Answer by orca(336)

To find the value of x, we need to solve the following equation:
sqrt(3x)+1=x+1
sqrt(3x)=x)
3x=x^2 (Square both sides)
x^2-3x=0
x(x-3)=0
So x = 0 or x = 3